OpenStudy (anonymous):

Please be 100% sure you know this answer before you answer it. I will fan and medal Jill began to solve an equation. Which expression below shows the next step Jill needs to take? JILLS SOLUTION __________________ / =divided by y/(-2) = 4.5 y/(-2)•(-2)=?

OpenStudy (anonymous):

@KoalaGurl1 @thread252

OpenStudy (thread252):

is this multiple choice?

OpenStudy (anonymous):

@thread252 nope

OpenStudy (thread252):

?????? that is bull crap! how are they gonna give us this kind of question and not answer choice???????????????

OpenStudy (anonymous):

@thread252 😂😂 it's ok

OpenStudy (anonymous):

@Ddos_Dragon @KoalaGurl1 @welshfella

OpenStudy (ddos_dragon):

Jill needs to multiply 4.5 by negative two.

OpenStudy (ddos_dragon):

Not 100% Sure, more 95% sure :P

OpenStudy (anonymous):

Do you know any other people that can help? @Ddos_Dragon ?

OpenStudy (ddos_dragon):

If y/(-2) = 4.5, then we can switch out y/(-2) with 4.5 4.5•(-2)=? Her next and final step is to multiply 4.5 by negative two. By the way the answer to the equation 4.5 * (-2) = -9

OpenStudy (anonymous):

Ok NEXY QUESTION!😃

OpenStudy (ddos_dragon):

If she is trying to find the variable that is wrong. Thats where the 5% doubt comes in :P

OpenStudy (ddos_dragon):

Was it correct?

OpenStudy (anonymous):

I don't know yet but I'm sure it was I will tell you at the end!!

OpenStudy (anonymous):

@Ddos_Dragon

OpenStudy (ddos_dragon):

Okay :)

OpenStudy (koalagurl1):

4.5*(-2) right?

OpenStudy (koalagurl1):

It's -9.

OpenStudy (anonymous):

Yeah I think so @KoalaGurl1

OpenStudy (koalagurl1):

Because 4.5*2 = 9 but its negative so its -9

OpenStudy (koalagurl1):

oops i guess @Ddos_Dragon already explained, didn't see that sorry. :)